### v3

@@ Abstract
In this paper, we prove that it is always possible to limit the number of consecutive degenerate pivots that the simplex algorithm performs to $n-m-1$, where $n$ is the number of variables and $m$ is the number of equality constraints of a given LP in standard equality form.
Keywords:
Simplex Algorithm Linear Programming Degeneracy.
@@ 1 Introduction
The simplex algorithm relies on the concept of basis, where a basis corresponds to an inclusion-wise minimal set of tight constraints which defines an extreme point solution of the underlying linear program. In each step, it performs a basis exchange by replacing one constraint currently in the basis with a different one. Such an exchange is called a pivot. From a geometric perspective, a basis exchange identifies a direction to move from the current extreme point. The simplex algorithm only considers pivots that yield an improving direction, with respect to the objective function to be optimized. As the result of pivoting, it is possible for the algorithm to either stay at the same extreme point, or to move to an adjacent extreme point of the underlying polyhedron. We refer to a pivot of the former type as degenerate, in contrast to a pivot of the latter type, which will be called non-degenerate. The pivot rule which determines how to perform these basis exchanges (i.e., deciding which inequality enters and which inequality leaves the basis) is the core of the simplex algorithm. Various pivot rules have been introduced over decades, but none of them yields a polynomial time version of the simplex algorithm so far (see, e.g., [3, 14, 15, 16, 18, 38, 23, 44] and the references therein).
@@
When analyzing the performance of the simplex algorithm, one can often assume that the underlying polyhedron $\mathcal{P}$ is non degenerate, e.g., by performing a small perturbation of the constraints right-hand-side. In this case, a necessary condition to guarantee an efficient performance of the simplex algorithm, is that there is always a way to reach an optimal solution within a polynomial number of non-degenerate pivots. Geometrically, this would imply that there is always a path between two extreme points on the 1-skeleton of $\mathcal{P}$ (the graph where vertices correspond to the extreme points of $\mathcal{P}$, and edges correspond to the 1-dimensional faces of $\mathcal{P}$), whose length is bounded by a polynomial in the input description of $\mathcal{P}$. This is a major and still open question in the field (see for example [6, 12, 25, 34, 40]).
@@
On the other hand, even in the case when a short path exists, it is not clear how the simplex algorithm would be bound to follow it (see for instance the case of $0/1$-polytopes discussed in [4]). In fact, given any LP one can apply some transformation to it to ensure the existence of a short path between an initial extreme point solution and an optimal one in the underlying polyhedron (such as, e.g., relying on an extended formulation with one extra dimension to construct a pyramid whose apex is adjacent to all extreme points of the initial polyhedron). Such standard transformations usually make the polyhedron become highly degenerate. In presence of degeneracy, an extreme point might have exponentially many distinct bases defining it, and the simplex algorithm might perform an exponential number of consecutive degenerate pivots, staying in the same vertex, (see, e.g., [10]). Such behavior is referred to as stalling. In some pathological cases, certain pivot rules might even provide an infinite sequence of consecutive degenerate pivots at a degenerate vertex, a phenomenon called cycling. Although cycling can be easily avoided by employing, for instance, Bland’s rule [5] or a lexicographic rule [39], there is no known pivot rule that prevents stalling. As stated in several papers (e.g., [27, 28]), it is well known that solving a general LP in strongly polynomial time can be reduced to finding a pivot rule that prevents stalling in general polyhedra. However, there are a few cases in which it is known that the issue of stalling can be handled. The most famous example is the class of transportation polytopes, for which pivot rules with polynomial bounds on the number of consecutive degenerate pivots were introduced in [10] and further developed in [1, 17, 32]. See also the work of [31] for a strongly polynomial version of the primal network simplex.
@@
Our findings are inspired by [19] to a great extent. The latter work shows that, given an LP in standard equality form with a totally unimodular coefficient matrix $A\in\mathbb{\{}-1,0,1\}^{m\times n}$, a vertex solution $x$ and an improving feasible circuit direction for $x$, one can construct a pivot rule which performs at most $m$ consecutive degenerate pivots. We generalize their proof, and show that one can employ any improving direction at a vertex $x$ of a general polyhedron, to avoid stalling. Our result, proved in Section 2, is summarized in the following theorem.
@@
Given any LP of the form $\min\{c^{T}x:Ax=b,x\geq 0\}$ with $A\in\mathrm{\hskip 0.28453ptIR}^{m\times n}$, there exists a pivot rule that limits the number of consecutive degenerate simplex pivots at any non-optimal extreme solution to at most $n-m-1$.
We stress here that the pivots considered in Theorem 1.1 are degenerate simplex pivots, meaning that each of them yields an improving direction at the given extreme point, though this direction is not feasible. It is important to point this out as in general, given two adjacent extreme points $x,x^{\prime}$, one can easily perform a sequence of basis exchanges that yield $x^{\prime}$ from $x$ by identifying the common tight linearly independent constraints, and introducing each of them to the current basis in any order, until the direction $x^{\prime}-x$ is seen. However, we here want a strategy that guarantees that each basis exchange is realizable by the simplex algorithm and hence defines an improving direction.
@@ 2.1 Preliminaries
The simplex algorithm works with LP in standard equality form:
$\begin{array}[]{rl}\min&c^{T}x\\ s.t.&Ax=b\\ &x\geq 0\end{array}$ (1)
@@
For $A\in\mathrm{\hskip 0.28453ptIR}^{m\times n}$ and $B\subseteq[n]$ we use $A_{B}$ to denote the submatrix of $A$ formed by the columns indexed by $B$. Similarly, for $x\in\mathrm{\hskip 0.28453ptIR}^{n}$ and $B\in[n]$, the notation $x_{B}$ is used to denote a vector consisting of the entries of $x$ indexed by $B$. A basis of (1) is a subset $B\subseteq[n]$ with $|B|=m$ and $A_{B}$ being non-singular. The point $x$ with $x_{B}=A_{B}^{-1}b$, $x_{N}=\mathbb{0}$ where $N:=[n]\setminus B$ is a basic solution of (1) with basis $B$. If additionally $x_{B}\geq 0$, both the basic solution $x$ and the basis $B$ are feasible. If $x_{i}>0$ for all $i\in B$, then $B$ and $x$ are called non-degenerate and degenerate otherwise, i.e., if $x_{i}=0$ for some $i\in B$. We let $\overline{A}:=A_{B}^{-1}A_{N}$ and $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$. In particular, $\overline{c}_{N}\in\mathrm{\hskip 0.28453ptIR}^{N}$ is the vector of so-called reduced costs for the basis $B$. The coordinates of the reduced cost vector will be addressed as $\overline{c}_{N,i}$, where the first subscript will be dropped if the basis $B=[n]\setminus N$ is clear from the context.
@@
The simplex algorithm considers in each iteration a feasible $B$. If all elements of $\overline{c}_{N}$ are non-negative, then the basis $B$ and the corresponding basic feasible solution $x$ are known to be optimal. Otherwise, the algorithm pivots by choosing a non-basic coordinate with negative reduced cost to enter the basis, say $f$. It then performs a minimum ratio test to compute an index $i$ that minimizes $\frac{x_{i}}{\overline{A}_{if}}$ among all indices $i\in B$ for which $\overline{A}_{if}>0$. Such index $i$ will be the one leaving the basis, and it corresponds to a basic coordinate which hits its bound first when moving along the direction given by the tight constraints indexed by $B\setminus\{f\}$. At each iteration there could be multiple candidate indices for entering the basis (all the ones with negative reduced cost), as well as multiple candidate indices to leave the basis (all the ones for which the minimum ratio test value is attained). The choice of the entering and the leaving coordinates is the essence of a pivot rule (see [38] for a survey). A pivot is degenerate if the attained minimum ratio test value is 0 (hence, the extreme point solution $x$ does not change), and non-degenerate otherwise. Note that only coordinates with negative reduced cost are considered for entering the basis, since only such a choice guarantees the simplex algorithm to make progress when a non-degenerate pivot occurs. To later emphasize that we only refer to that kind of improving pivots, we will call them simplex pivots.
@@
Finally, we let $\mathrm{kern}(A):=\{y\in\mathrm{\hskip 0.28453ptIR}^{n}:Ay=0\}$. Given (1), we call a vector $y\in\mathrm{kern}(A)$ with $c^{T}y<0$ an improving direction. Such $y$ is said to be feasible at a basic feasible solution $x$ if $x+\epsilon y\geq 0$ for sufficiently small positive $\epsilon$. Note that for any $y\in\mathrm{kern}(A)$ and any basis $B$ of $A$, the following holds:
@@ 2.2 A geometric intuition
[…]
@@
It is well known, that a basic feasible solution $x$ of (1) with basis $B$ is not optimal if and only if there exists an improving feasible direction, i.e., there exists $y^{0}\in\mathrm{\hskip 0.28453ptIR}^{n}$ such that $Ay^{0}=0$, $c^{T}y^{0}<0$ and $y^{0}_{i}\geq 0$ for all $i\in[k]\cup\big{(}[n]\setminus[m]\big{)}$. Consider any such $y^{0}$. Let $Q_{1}(y^{0},B):=\{i\in[k]\mid y^{0}_{i}>0\}$ and $Q_{2}(y^{0},B):=\{i\in[n]\setminus[m]\mid y^{0}_{i}>0\}$. Without loss of generality, let $Q_{1}(y^{0},B)=[r]$ with $0\leq r\leq k$ and $Q_{2}(y^{0},B)=\{m+1,\dots,m+t\}$ with $1\leq t\leq n-m$. See Table 1 for an illustration.
@@ 1 $r$ $k$ $m$ $f$ $m+t$ $n$
$z$star $\star$ $0_{+}$ $0_{+}$star $\star$ 1 0 0 0 0
$y^{0}+\alpha z$ 0 $0_{+}$ $0_{+}$ $0_{+}$ $0_{+}$star $\star$ + + + 0 0
Table 1: An illustration of entries of $x,y^{0},z$ and $y^{1}:=y^{0}+\alpha z$. A positive, a non-negative, a negative, and a sign-arbitrary entry is denoted by $+$, $0_{+}$, -- and $\star$, respectively. Without loss of generality, we here assumed that the entering index $f$ is $m+1$, and that an index $g$ for which the minimum in Case III is attained is 1.
@@
In particular, this means $Eq^{>}(f)\subseteq[m]\setminus[k]$. We perform a non-degenerate pivot, by selecting an index that minimizes ()(*)( ∗ ) as the leaving index.
@@
Case III: $\overline{A}_{if}\leq 0$ for all $i\in[k]\setminus[r]$ and $\overline{A}_{if}>0$ for some $i\in[r]$. In this case, we are going to change our improving feasible direction. Consider the following vector $z\in\mathrm{\hskip 0.28453ptIR}^{n}$ which is, in fact, an improving (thought not feasible) direction for the entering variable $x_{f}$:
$z_{i}=\begin{cases}-\overline{A}_{if},&\text{for $i\in[m]$}\\ 1,&\text{for $i=f$}\\ 0,&\text{otherwise}\end{cases}$ (3)
@@
$y^{1}:=y^{0}+\alpha z\;\mbox{ with }\;\alpha:=\min_{i\in Eq^{>}(f)\cap[r]}\frac{y_{i}^{0}}{|z_{i}|}>0\,.$
@@
The key observation is that when the third case of the above case distinction has occurred, repeating the same for $B$ and $y^{1}$ falls into the second case (because the basis $B$ has not changed, meanwhile $y^{1}_{g}=x_{g}=0$ with $g\in B$ and $\overline{A}_{gf}>0$). Therefore, after each degenerate pivot, the cardinality of the support of the improving feasible direction in the non-basic indices decreases. Hence, a sequence of $|Q_{2}(y^{0},B)|$ degenerate pivots would yield an improving feasible direction $y^{\prime}$ and a basis $B^{\prime}$ of $x$ with $Q_{2}(y^{\prime},B^{\prime})=0$, which in turn implies $0=\overline{c}^{T}_{N^{\prime}}y^{\prime}_{N^{\prime}}=c^{T}y^{\prime}$ due to (2), yielding a contradiction. Hence the number of consecutive degenerate pivots with the suggested pivot rule cannot exceed $n-m-1$. ∎
We conclude with a remark which will be useful in the next section.
Remark 1
Let $x$ be the currently considered basic feasible solution with basis $B$. If $y^{0}$ is chosen to be equal to $\tilde{x}-x$ for some basic feasible solution $\tilde{x}$ of our LP with $c^{T}\tilde{x}>c^{T}x$, then one can observe that $|Q_{2}(y^{0},B)|\leq m$ because $\tilde{x}$ has at most $m$ non-zero coordinates. By the proof of Theorem 1.1 above, the latter value strictly decreases with each degenerate step produced by the antistalling pivot rule. Hence, the total number of consecutive degenerate pivots at the vertex $x$ can be strengthen to be at most $\min\{n-m-1,m-1\}$.

### v4

@@ Abstract
In this paper, we prove that one can employ any improving feasible direction at an extreme point to limit the number of consecutive degenerate pivots that the simplex algorithm performs to nm1n-m-1, where $n$ is the number of variables and $m$ is the number of equality constraints of a given LP in standard equality form.
@@ 1 Introduction
The simplex algorithm relies on the concept of basis, where a basis corresponds to an inclusion-wise minimal set of tight constraints which defines an extreme point solution of the underlying linear program. In each step, it performs a basis exchange by replacing one constraint currently in the basis with a different one that is not in the basis. Such an exchange is called a pivot. From a geometric perspective, a basis exchange identifies a direction to move from the current extreme point. The simplex algorithm only considers pivots that yield an improving direction, with respect to the objective function to be optimized. As the result of pivoting, it is possible for the algorithm to either stay at the same extreme point, or to move to an adjacent extreme point of the underlying polyhedron. We refer to a pivot of the former type as degenerate, in contrast to a pivot of the latter type, which will be called non-degenerate. The pivot rule which determines how to perform these basis exchanges (i.e., deciding which inequality enters and which inequality leaves the basis) is the core of the simplex algorithm. Various pivot rules have been introduced over decades, but none of them yields a polynomial time version of the simplex algorithm so far (see, e.g., [3, 16, 20, 18, 21, 41, 26, 47] and the references therein).
@@
When analyzing the performance of the simplex algorithm, one can often assume that the underlying polyhedron $\mathcal{P}$ is non degenerate, e.g., by performing a small perturbation of the constraint’s right-hand-side. In this case, a necessary condition to guarantee an efficient performance of the simplex algorithm, is that there is always a way to reach an optimal solution within a polynomial number of non-degenerate pivots. Geometrically, this would imply that there is always a path between two extreme points on the 1-skeleton of $\mathcal{P}$ (the graph where vertices correspond to the extreme points of $\mathcal{P}$, and edges correspond to the 1-dimensional faces of $\mathcal{P}$), whose length is bounded by a polynomial in the input description of $\mathcal{P}$. This is a major and still open question in the field, commonly referred to as the polynomial Hirsch Conjecture (see for example [6, 12, 27, 37, 43]).
@@
On the other hand, even in the case when a short path exists, it is not clear how the simplex algorithm would be bound to follow it (see for instance the case of $0/1$-polytopes discussed in [4], or the hardness results in [15, 9]). In fact, given any LP one can apply some transformation to it to ensure the existence of a short path between an initial extreme point solution and an optimal one in the underlying polyhedron (such as, e.g., relying on an extended formulation with one extra dimension to construct a pyramid whose apex is adjacent to all extreme points of the initial polyhedron). Such standard transformations usually make the polyhedron become highly degenerate. In the presence of degeneracy, an extreme point might have exponentially many distinct bases defining it, and the simplex algorithm might perform an exponential number of consecutive degenerate pivots, staying in the same vertex, (see, e.g., [11]). Such behavior is referred to as stalling. In some pathological cases, certain pivot rules might even provide an infinite sequence of consecutive degenerate pivots at a degenerate vertex, a phenomenon called cycling. Although cycling can be easily avoided by employing, for instance, Bland’s rule [5] or a lexicographic rule [42], there is no known pivot rule that prevents stalling. As stated in several papers (e.g., [28, 30]), it is well known that solving a general LP in strongly polynomial time can be reduced to finding a pivot rule that prevents stalling in general polyhedra, and that can be implemented in strongly polynomial time. However, there are a few cases in which it is known that the issue of stalling can be handled. The most famous example is the class of transportation polytopes, for which pivot rules with polynomial bounds on the number of consecutive degenerate pivots were introduced in [11] and further developed in [1, 19, 35]. See also the work of [33] for a strongly polynomial version of the primal network simplex.
@@
Our findings are inspired by [22] to a great extent. The latter work shows that, given an LP in standard equality form with a totally unimodular coefficient matrix $A\in\mathbb{\{}-1,0,1\}^{m\times n}$, a vertex solution $x$ and an improving feasible circuit direction for $x$, one can construct a pivot rule which performs at most $m$ consecutive degenerate pivots. We generalize their proof, and show that one can employ any improving feasible direction at a vertex $x$ of a general polyhedron, to avoid stalling. Our result, proved in Section 2, is summarized in the following theorem.
@@
Given any LP of the form $\min\{c^{T}x:Ax=b,x\geq 0\}$ with $A\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{m\times n}$, there exists a pivot rule that limits the number of consecutive degenerate simplex pivots at any non-optimal extreme solution to at most nm1n-m-1.
The pivot rule can be efficiently implemented whenever an improving feasible direction at a non-optimal extreme point is available.
We stress here that the pivots considered in Theorem 1.1 are degenerate simplex pivots, meaning that each of them yields an improving direction at the given extreme point, though this direction is not feasible. It is important to point this out as in general, given two adjacent extreme points $x,x^{\prime}$, one can easily perform a sequence of basis exchanges that yields $x^{\prime}$ from $x$ by identifying the common tight linearly independent constraints, and introducing each of them to the current basis in any order, until the direction $x^{\prime}-x$ is seen. However, we here want a strategy that guarantees that each basis exchange defines an improving direction, and hence is realizable by the simplex algorithm.
@@ 2.1 Preliminaries
The simplex algorithm works with LPs in standard equality form:
$\begin{array}[]{rll}\min&c^{T}x&\\ \mbox{s.t.}&Ax&=b\\ &x&\geq 0\end{array}$ (1)
@@
For $A\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{m\times n}$ and $B\subseteq[n]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=\{1,2,\dots,n\}}$, we use $A_{B}$ to denote the submatrix of $A$ formed by the columns indexed by $B$. Similarly, for $x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}$ and $B\in[n]$, the notation $x_{B}$ is used to denote a vector consisting of the entries of $x$ indexed by $B$. A basis of (1) is a subset $B\subseteq[n]$ with $|B|=m$ and $A_{B}$ being non-singular. The point $x$ with $x_{B}=A_{B}^{-1}b$, $x_{N}=0$ where $N:=[n]\setminus B$ is a basic solution of (1) with basis $B$. If additionally $x_{B}\geq 0$, both the basic solution $x$ and the basis $B$ are feasible. If $x_{i}>0$ for all $i\in B$, then $B$ and $x$ are called non-degenerate. If instead $x_{i}=0$ for some $i\in B$, then $B$ and $x$ are called degenerate. We let $\overline{A}:=A_{B}^{-1}A_{N}$ and $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$. In particular, $\overline{c}_{N}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{N}$ is the vector of so-called reduced costs for the basis $B$. The coordinates of the reduced cost vector will be addressed as $\overline{c}_{N,i}$, where the first subscript will be dropped if the basis $B=[n]\setminus N$ is clear from the context.
@@
The simplex algorithm considers in each iteration a feasible $B$. If all elements of $\overline{c}_{N}$ are non-negative, then the basis $B$ and the corresponding basic feasible solution $x$ are known to be optimal. Otherwise, the algorithm pivots by choosing a non-basic coordinate with negative reduced cost to enter the basis, say $f$. It then performs a minimum ratio test to compute an index $i$ that minimizes $\frac{x_{i}}{\overline{A}_{if}}$ among all indices $i\in B$ for which $\overline{A}_{if}>0$. Such index $i$ will be the one leaving the basis, and it corresponds to a basic coordinate which hits its non-negativity bound first when moving along the direction given by the tight constraints indexed by $B\setminus\{f\}$. At each iteration there could be multiple candidate indices for entering the basis (all the ones with negative reduced cost), as well as multiple candidate indices to leave the basis (all the ones for which the minimum ratio test value is attained). The choice of the entering and the leaving coordinates is the essence of a pivot rule (see [41] for a survey). A pivot is degenerate if the attained minimum ratio test value is 0 (hence, the extreme point solution $x$ does not change), and non-degenerate otherwise. Note that only coordinates with negative reduced cost are considered for entering the basis, since only such a choice guarantees the simplex algorithm to make progress when a non-degenerate pivot occurs. To later emphasize that we only refer to that kind of improving pivots, we will call them simplex pivots.
@@
Finally, we let $\mathrm{kern}(A):=\{y\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}:Ay=0\}$. Given (1), we call a vector $y\in\mathrm{kern}(A)$ with $c^{T}y<0$ an improving direction. Such $y$ is said to be feasible at a basic feasible solution $x$ if $x+\varepsilon y\geq 0$ for sufficiently small positive $\varepsilon$. Note that for any $y\in\mathrm{kern}(A)$ and any basis $B$ of $A$, the following holds:
@@ 2.2 A geometric intuition
Before providing a formal proof, we give a geometric intuition on how our pivot rule works. For this, it will be easier to abandon the standard equality form. In particular, consider a degenerate vertex $x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d}$ of a polytope $\mathcal{P}$ of dimension $d\in\mathbbm{N}$ defined by inequalities $a_{i}^{T}x\leq b_{i}$ with $a_{i}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d},b_{i}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR},i\in[n]$. In this setting, degeneracy means that more than $d$ inequalities are tight at $x$. Consider a subset $N$ of the set of indices of all inequalities that are tight at $x$ with $|N|=d$ and let $B:=[n]\setminus N$ (see Figure 1). Note that here we intentionally redefine a notation used in the standard equality form, where non basic coordinates always have their corresponding constraints tight. Since $x$ is degenerate, there is at least one inequality that is tight at $x$ with index in $B$. We denote the subset of such inequalities by $W\subseteq B$. Finally, assume $x$ is not an optimal vertex of $\mathcal{P}$ when minimizing an objective function $c\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d}$ over $\mathcal{P}$, and let $y^{0}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{d}$ be an improving feasible direction at $x$, i.e., such that $x+\varepsilon y^{0}\in\mathcal{P}$ for a sufficiently small positive $\varepsilon$ and $c^{T}y^{0}<0$. In order to find an improving edge of $\mathcal{P}$ at $x$, we look at the directions given by the extremal rays of the basic cone $C(N):=\{x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}\mid a_{i}^{T}x\leq 0,i\in N\}$. If there is an improving feasible direction among them, we are done. Otherwise, we reduce the dimension of the polytope in the following way. We pick an extremal ray $z\in C(N)$ such that $c^{T}z<0$. Note that $z$ is formed by d1d-1 inequalities from $N$. Let $f\in N$ be the only inequality index not used to define $z$, among the $d$ tight inequalities in $C(N)$ ($f$ is depicted as the gray fading facet in Figure 1). Consider a vector combination $y^{0}+\alpha z$ where $\alpha\geq 0$ and note that it provides an improving direction for any positive $\alpha$. Since $y^{0}$ is contained in the feasible cone $C(N\cup W)=\{x\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}\mid a_{i}^{T}x\leq 0,i\in N\cup W\}$ but $z$ is not, the vector $y^{0}+\alpha z$ leaves $C(N\cup W)$ for sufficiently large $\alpha$. Hence there has to be a number $\alpha^{1}\geq 0$ such that $y^{0}+\alpha^{1}z$ is contained in a facet of $C(N\cup W)$. Without loss of generality assume that the latter facet is defined by the $g\textsuperscript{th}$ inequality (green facet in Figure 1). Note that $g\in W$ since $y^{0}$ and $z$ are both contained in the basic cone $C(N)$ and so is $y^{0}+\alpha z$ for any non-negative $\alpha$. Then, we define a new improving feasible direction $y^{1}:=y^{0}+\alpha^{1}z$, $N=\big(N\setminus\{f\}\big)\cup\{g\}$, $B=\big(B\setminus\{g\}\big)\cup\{f\}$ and continue searching for an improving edge inside the facet defined by the $g\textsuperscript{th}$ inequality. Since $x+\varepsilon y^{1}$ belongs to this facet for an $\varepsilon>0$ small enough, the latter yields dimension reduction. After at most d1d-1 such steps we are bound to encounter a one-dimensional face of $\mathcal{P}$ containing $x+\varepsilon y^{\prime}$, where $y^{\prime}$ is an improving feasible direction at $x$. For an illustration see Figure 1.
Figure 1: To the left, facets corresponding to inequalities in NN and $W$ are colored in navy and white, respectively. To the right, the gray fading facet is defined by the $f\textsuperscript{th}$ inequality and the green one corresponds to the $g\textsuperscript{th}$.
@@
It is well known that a basic feasible solution $x$ of (1) with basis $B$ is not optimal if and only if there exists an improving feasible direction, i.e., there exists $y^{0}\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}$ such that $Ay^{0}=0$, $c^{T}y^{0}<0$ and $y^{0}_{i}\geq 0$ for all $i\in[k]\cup\big([n]\setminus[m]\big)$. Consider any such $y^{0}$. Let $Q_{1}(y^{0},B):=\{i\in[k]\mid y^{0}_{i}>0\}$ and $Q_{2}(y^{0},B):=\{i\in[n]\setminus[m]\mid y^{0}_{i}>0\}$. Without loss of generality, let $Q_{1}(y^{0},B)=[r]$ with $0\leq r\leq k$ and $Q_{2}(y^{0},B)=\{m+1,\dots,m+t\}$ with $1\leq t\leq n-m$. See Table 1 for an illustration.
@@ 1 $r$ $k$ $m$ $f$ $m+t$ $n$
$z$ $\star$ $\star$ $_{+} 0+0_{+}$ $\star$ $\star$ 1 0 0 0 0
$y^{0}+\alpha z$ 0 $_{+} 0+0_{+}$ $_{+} 0+0_{+}$ $\star$ $\star$ + + + 0 0
Table 1: An illustration of entries of $x,y^{0},z$ and $y^{1}:=y^{0}+\alpha z$. A positive, a non-negative, a negative, and a sign-arbitrary entry is denoted by ++, $0_{+}$, - and $\star$, respectively. Without loss of generality, we here assumed that the entering index $f$ is $m+1$, and that an index $g$ for which the minimum in Case III is attained is 1.
@@
In particular, this means $Eq^{>}(f)\subseteq[m]\setminus[k]$. We perform a non-degenerate pivot, by selecting an index that minimizes ()(*) as the leaving index.
@@
Case III: $\overline{A}_{if}\leq 0$ for all $i\in[k]\setminus[r]$ and $\overline{A}_{if}>0$ for some $i\in[r]$. In this case, we are going to change our improving feasible direction. Consider the following vector $z\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{n}$ which is, in fact, an improving (though not feasible) direction for the entering variable $x_{f}$:
$z_{i}=\begin{cases}-\overline{A}_{if}&\text{for $i\in[m]$,}\\ 1&\text{for $i=f$,}\\ 0&\text{otherwise.}\end{cases}$ (3)
@@
$y^{1}:=y^{0}+\alpha z\;\mbox{ with }\;\alpha:=\min_{i\in Eq^{>}(f)\cap[r]}\frac{y_{i}^{0}}{-z_{i}}>0\,.$
@@
The key observation is that when the third case of the above case distinction has occurred, repeating the same for $B$ and $y^{1}$ falls into the second case (because the basis $B$ has not changed, meanwhile $y^{1}_{g}=x_{g}=0$ with $g\in B$ and $\overline{A}_{gf}>0$). Therefore, after each degenerate pivot, the cardinality of the support of the improving feasible direction in the non-basic indices decreases. Hence, a sequence of $|Q_{2}(y^{0},B)|$ degenerate pivots would yield an improving feasible direction $y^{\prime}$ and a basis $B^{\prime}$ of $x$ with $Q_{2}(y^{\prime},B^{\prime})=0$, which in turn implies $0=\overline{c}^{T}_{N^{\prime}}y^{\prime}_{N^{\prime}}=c^{T}y^{\prime}$ due to (2), yielding a contradiction. Hence the number of consecutive degenerate pivots with the suggested pivot rule cannot exceed nm1n-m-1. ∎
For the sake of clarity, the antistalling pivot rule is formally summarized in Algorithm 1.
Algorithm 1 Antistalling pivot rule.
0: A basic feasible solution $x$ of (1) with basis $B=\{1,\dots,m\}$ and an improving feasible direction $y$.
1: Compute the matrix $\overline{A}:=A_{B}^{-1}A_{N}$ and the reduced cost vector $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$, where $N:=[n]\setminus B$.
2: Compute the entering index $f$ as $\mathrm{argmin}_{\{i\in N:\overline{c}_{i}<0\}\cap\mathrm{supp}(y)}\overline{c}_{i}$.
3:if $\overline{A}_{if}\leq 0$ then
4:  The optimal cost is unbounded: the algorithm terminates.
5:else
6:  Perform the minimum ratio test:
$r:=\min_{i\in\mathrm{supp}\big(\overline{A}_{if}\big)}\frac{x_{i}}{\overline{A}_{if}}$ and $J:=\mathrm{argmin}_{i\in\mathrm{supp}\big(\overline{A}_{if}\big)}\frac{x_{i}}{\overline{A}_{if}}$.
7:  if $r>0$ (case I) then
8:   Choose $j\in J$ as the leaving variable. Perform non-degenerate pivot: $B=B\cup\{f\}\setminus\{j\}$.
9:  else if $J\cap B\setminus\mathrm{supp}(y)\neq\emptyset$ (case II) then
10:   Choose $j\in J\cap B\setminus\mathrm{supp}(y)$ as the leaving variable. Perform degenerate pivot: $B=B\cup\{f\}\setminus\{j\}$. Go back to 1.
11:  else
12:   (case III) Compute improving (though not feasible) direction $z$ as:
$z_{i}=\begin{cases}-\overline{A}_{if}&\text{for $i\in B$,}\\ 1&\text{for $i=f$,}\\ 0&\text{otherwise.}\end{cases}$
Compute the new improving feasible direction $y:=y+\alpha z$, where $\alpha:=\min_{i\in\mathrm{supp}(\overline{A}_{if})\cap\mathrm{supp}(y)\setminus\mathrm{supp}(x)}\frac{y_{i}}{\overline{A}_{if}}$. Go back to 3.
13:  end if
14:end if
We conclude with a strengthening of the bound of Theorem 1.1 which will be useful in the next section.
Theorem 2.1
Given any LP of the form $\min\{c^{T}x:Ax=b,x\geq 0\}$ with $A\in\mathrm{\hskip 0.28453ptI\hskip-2.27621ptR}^{m\times n}$, there exists a pivot rule that limits the number of consecutive degenerate simplex pivots at any non-optimal extreme solution $x$ to at most $\min\{n-m-1,m-1\}$.
The pivot rule can be efficiently implemented whenever an improving feasible direction of the form $\tilde{x}-x$, for some extreme point solution $\tilde{x}$, is available.
Proof
Let $x$ be the currently considered basic feasible solution with basis $B$. If $y^{0}$ is chosen to be equal to $\tilde{x}-x$ for some basic feasible solution $\tilde{x}$ of our LP with $c^{T}\tilde{x}<c^{T}x$, then one can observe that $|Q_{2}(y^{0},B)|\leq m$ because $\tilde{x}$ has at most $m$ non-zero coordinates. By the proof of Theorem 1.1 above, the latter value strictly decreases with each degenerate step produced by the antistalling pivot rule. Hence, the total number of consecutive degenerate pivots at the vertex $x$ can be strengthened to be at most $\min\{n-m-1,m-1\}$. ∎